MSG Corporation issued $102,000 of 3-year, 5% bonds outstanding on December 31, 2020 for $107,000. The bonds pay interest annually and MSG uses straight-line amortization. On May 1, 2021, $10,200 of the bonds were retired at 112. As a result of the retirement, MSG will report: (Do not round intermediate calculations and round final answer to nearest whole dollar.)

Multiple Choice

  • a $724 loss.

  • a $780 loss.

  • a $1,448 loss.

  • a $1,448 gain.
Answer #1
Issue price of $10,200 bonds 10700 =107000/102000*10200
Premium on issue 500 =10700-10200
Premium amortization for 4 months 56 =500/3*4/12
Carrying value on May 1 10644 =10700-56
Bond redemption price 11424 =10200*1.12
Less: Carrying value on May 1 10644
Loss on redemption 780 loss
Option B a $780 loss is correct option